Use a map scale of 2 inches:25 miles fo find the actual distance for each measurment.
1.) 3 in.=
2.) 1/2 in.=
3.) 5 in.=
4.Anne lives 30 miles from her cousin.On a map she measures 3 in.between their homes.What is the scale of the map?

Answer by algebrapro18(249)

1) 2/25 = 3/x
2x=75
x = 37.5 miles

2) 2/25 = .5/x
2x=25/2
x = 25/4
x=6.25 miles

3)2/25 = 5/x
2x=125
x = 62.5 miles

well there are 3in/30miles so 1in/10 miles is the scale of the map.